The driver of a speeding empty truck slams on the brakes and slows down to a stop through a distance d. On a second trial, the truck carries a load that doubles its mass. What distance will the loaded truck cover to come to rest, when it is moving at the same speed and the brakes are slammed?
1
Expert's answer
2020-10-13T07:08:42-0400

Solution

Friction force is given by

"F=\mu N=\mu mg"

where N is normal force.

Using energy conservation

"Fd=\Delta KE=\frac{mv^2}{2}"

"d=\frac{mv^2}{2F}=\frac{mv^2}{2\mu mg}"

"d=\frac{v^2}{2\mu g}"

This equation clearly shows that distance d does not depend on mass.

Therefore d remains same.